Shift report 20000807 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Aug 07 2000 - 00:02:22 EDT

  • Next message: http account: "Shift report 20000807 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Aug  7 00:02:21 2000, Shift: 16:00-24:00
    Supervisor: YB, with: TT, JN
    Log book#4 pages 14 - 20
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Cogged beam at gamma = 70 for over than 3 hours 
      (since about 8:40pm)
    2. - Runs 2076, 2079 and 2080 still running. Very low rate.
         run 2079 has about 200 collisions and about 700 in 
         run2080 so far with D5 on.
       - Trigger's timing was fine-tuned by FV. ZDC times are down 
         to one peak and BB has two timing peaks.
       - KO introduced "HV health monitor" at:
         http://pii3.brahms.bnl.gov/~daq/opus-home-daq-run/hvStatus.html
         Very helpfull!
    3. - FFS1440 (H1 and C1) is still down
       - BBL HV is tripping all the time for different channels,
         but mainly Ch.5. Seems like a mainframe problem.
       - Vertex is still off by about 8-10nsec (agrees by BB and 
         ZDC)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Aug 07 2000 - 00:06:08 EDT